### v1

@@ III RBM-based emulations
This section centers around the RBM emulators used to interpolate, extrapolate, or emulate $\mathcal{A}$ in a combined space of complex $E$ plane and the $\bm{\theta}$ dimensions. Note some parameters inside $\bm{\theta}$ could be shared among the $H$ operator and the two sources. The training points, labeled as $(E^{tr}_{\alpha},\bm{\theta}^{tr}_{\alpha})$ with $\alpha=1,,,N_{b}$, are those with positive $\mathrm{Im}\,E^{tr}$. With those parameter values or at those points in the parameter space, we solve Eqs. (3) and (4) to get the training solution snapshots, $|\Psi^{tr}_{\alpha}\rangle\equiv|\Psi(E_{\alpha}^{tr},\bm{\theta}_{\alpha}^{tr})\rangle$ and $\langle\tilde{\Psi}^{tr}_{\alpha}|\equiv\langle\tilde{\Psi}(E_{\alpha}^{tr},\bm{\theta}_{\alpha}^{tr})|$. This step is called the training or off-line step. In the next emulation or online step, the RBM proposes a general solution at emulation point $(E,\bm{\theta})$,
$|\Psi(E,\bm{\theta})\rangle$ $=\sum_{\alpha=1}^{N_{b}}c_{\alpha}(E,\bm{\theta})|\Psi_{\alpha}^{tr}\rangle\ .$ (8)
@@
As shown later, the inhomogeneous Schrödinger equations are effectively projected into a subspace spanned by the training solutions, $|\Psi^{tr}_{\alpha}\rangle$ and $\langle\tilde{\Psi}^{tr}_{\alpha}|$.
@@
As shown in previous studies, the RBM-based emulators have robust extrapolation capabilities because the emulators learn a problem’s solution manifold—mapped out by varying the input parameters—by sampling solution vectors in the subspace Duguet et al. (2024). Using our CEE, we extrapolate $\mathcal{A}$ at complex training energies to the real energy axis in the complex $E$ plane and further into the region where potential resonances could exist, i.e., this procedure extracts $\mathcal{A}$ at real energies and resonance properties from the training solutions, which satisfy bound-state-like boundary conditions Efros (1985b). In addition, the CERPE emulates those continuum physics extractions in the $\bm{\theta}$ space.
@@
When emulating at real energies with $\mathrm{Im}\,(E)\to 0^{+}$, the constructed wave function in Eq. (8) has bound-state-like asymptotic behavior, which is incorrect and should be outgoing waves oscillating to infinity spatially (see Fig. 2 for a visual demonstration). How is it possible for the emulated $\mathcal{A}$ at real energies to be sensible? The answer comes from the fact that the sources are spatially localized; $\mathcal{A}(E)$ can thus be determined by the overlap of the source with the internal part of the emulated solution $|\Psi\rangle$ as shown in Eq. (5). This basic setup allows us to emulate continuum physics, including $\mathcal{A}(E)$ at real $E$ and compressed spectra (bound, resonance, and DC states), without reproducing the intricate wave function asymptotics correctly. This point will be illustrated further using an example in Fig. 2 and the discussion around it.
@@
From the discussion at the end of Sec. II.3, we can see that $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s discretized BCTs (equivalently, the DC states in the compressed spectrum) are responsible for $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s near-threshold behavior. As shown in the Secs. IV and V, these poles exponentially cluster toward the branch points or thresholds, similar to the pole distribution in the (near)-optimal rational approximation. The threshold locations depend only on a particular $H$, but will the distribution of the discretized BCT poles be sensitive to the sources, i.e., different dynamics? The numeric results in the Secs. IV and V suggest the answer is no, which is consistent with the distribution being the property of a spectrum (albeit compressed) but not a property unique to specific dynamics. This is also in parallel with the observation that the “widespread” behavior of the pole distribution in the (near)-optimal rational approximation of functions with branch points is determined mainly by the locations of the branch points but not significantly by the other details of the functions Trefethen et al. (2021); Trefethen (2023).
@@
Moreover, the exponential clustering of the discretized BCT poles also means we need a small number of them (i.e., low-dim subspace) to approximate the smooth component of $\mathcal{A}$—defined as the full $\mathcal{A}$ with the resonance contributions subtracted—over an extensive range of real energies (see e.g., Fig. 4). This explains the feasibility of compressing the DC component of the continuum spectra. Of course, the bound and resonance states are the other spectra components, but their number can not be reduced.
@@
Now, the task is to determine $c_{\alpha}(E,\bm{\theta})$’s dependence on $E$ and $\bm{\theta}$, as addressed below. As a reminder, if we fix $\bm{\theta}$ to a specific value without varying them, the emulation is called CEE, and otherwise, it is named CERPE.
@@ III.1 A variational method for solving a linear system and the associated RBM ROM
Suppose we need to solve an inhomogeneous linear equation:
@@
and estimate $\langle\tilde{S}|\Psi\rangle$. A variational approach exists for this task Pomraning (1965). The involved functional, depending on both trial solution $|\Psi_{t}\rangle$ and an auxiliary variable $|\tilde{\Psi}_{t}\rangle$, is expressed as
@@
Suppose a trial solution pair is close to the exact solutions, i.e., $|\Psi_{t}\rangle=|\Psi\rangle+|\delta\rangle$ and $|\tilde{\Psi}_{t}\rangle=|\tilde{\Psi}\rangle+|\tilde{\delta}\rangle$, with $|\Psi\rangle$ as Eq. (9)’s exact solution and $|\tilde{\Psi}\rangle$ as the exact solution of another equation to be defined, we then have
@@
Therefore, we require $|\tilde{\Psi}\rangle$ satisfy
$\langle\tilde{S}|$ $=\langle\tilde{\Psi}|M\leftrightarrow|\tilde{S}\rangle=M^{\dagger}|\tilde{\Psi}\rangle\ ,$ (12)
so that the functional $\mathcal{F}$ is stationary around the exact solution and provides an estimate of $\langle\tilde{S}|\Psi\rangle$ on the second order of $|\delta\rangle\sim|\tilde{\delta}\rangle$. This equation defines the adjoint state vector $|\tilde{\Psi}\rangle$, which is why $|\tilde{\Psi}\rangle$ and Eq. (4) were introduced in Sec. I.
We will use the $\mathcal{F}$ functional to determine $c_{\alpha}$ as functions of $E$ and $\bm{\theta}$ in Eq. (8). This function guides us in choosing test functions in the framework of Galerkin projection. To construct a ROM for Eq. (9) with parameters in $M$ and the sources, we construct [^13] In the previous report Zhang (2024), we use the convention of the implicit sum over the repeated indices. Here, we express that sum explicitly.
$|\Psi_{t}\rangle$ $=\sum_{\alpha}c_{\alpha}|\Psi_{\alpha}^{tr}\rangle\ ,$ (13)
$\langle\tilde{\Psi}_{t}|$ $=\sum_{\alpha}\tilde{c}_{\alpha}\langle\tilde{\Psi}_{\alpha}^{tr}|\ ,$ (14)
@@
$M_{\alpha}|\Psi_{\alpha}^{tr}\rangle$ $=|{S}_{\alpha}\rangle\ ,$ (15)
$\langle\tilde{\Psi}_{\alpha}^{tr}|M_{\alpha}$ $=\langle\tilde{{S}}_{\alpha}|\leftrightarrow M_{\alpha}^{\dagger}|\tilde{\Psi}^{tr}_{\alpha}\rangle=|\tilde{{S}}_{\alpha}\rangle\ .$ (16)
@@
$\mathcal{F}(c,\tilde{c})$ $=\sum_{\alpha}\bigg{(}c_{\alpha}\langle\tilde{{S}}|\Psi_{\alpha}\rangle+\tilde{c}_{\alpha}\langle\tilde{\Psi}_{\alpha}|{S}\rangle\bigg{)}$
$\quad-\sum_{\alpha,\beta}\tilde{c}_{\alpha}c_{\beta}\langle\tilde{\Psi}_{\alpha}|M|\Psi_{\beta}\rangle\ .$ (17)
Note when ${S}$, $\tilde{{S}}$, or $M$ is without $\alpha.\beta$ indices, they are evaluated at the general emulation point. To find the stationary points, we look for the solutions such that the first derivatives of $\mathcal{F}$ with respect to $c$ and $\tilde{c}$ are zero. We then have a small-dimensional linear equation system:
$\begin{pmatrix}0&\begin{bmatrix}M\end{bmatrix}^{T}\\ \begin{bmatrix}M\end{bmatrix}&0\end{pmatrix}\begin{pmatrix}c\\ \tilde{c}\end{pmatrix}=\begin{pmatrix}\begin{bmatrix}\tilde{S}^{\dagger}\end{bmatrix}\\ \begin{bmatrix}S\end{bmatrix}\end{pmatrix}\ .$ (18)
The upper right corner of the matrix on the left side uses transpose, T, not complex conjugate, due to our definition of $\tilde{c}$ in Eq. (13). The block matrix,
$\begin{bmatrix}M\end{bmatrix}_{\alpha,\beta}\equiv\langle\tilde{\Psi}_{\alpha}|M|\Psi_{\beta}\rangle$ (19)
and similarly,
$\begin{bmatrix}\tilde{{S}}^{\dagger}\end{bmatrix}_{\alpha}$ $\equiv\langle\tilde{{S}}|\Psi_{\alpha}\rangle$ (20)
$\begin{bmatrix}{S}\end{bmatrix}_{\alpha}$ $\equiv\langle\tilde{\Psi}_{\alpha}|{S}\rangle\ .$ (21)
@@
$\sum_{\alpha}c_{\alpha}\langle\tilde{{S}}|\Psi_{\alpha}\rangle$ $=\sum_{\alpha}\tilde{c}_{\alpha}\langle\tilde{\Psi}_{\alpha}|{S}\rangle$
$=\sum_{\alpha,\beta}\tilde{c}_{\alpha}c_{\beta}\langle\tilde{\Psi}_{\alpha}|M|\Psi_{\beta}\rangle\ .$ (22)
We thus have
$\mathcal{F}|_{\mathrm{stationary}}$ $=\sum_{\alpha}c_{\alpha}\langle\tilde{S}|\Psi_{\alpha}\rangle$ (23)
Therefore, after obtaining, $|\Psi^{tr}_{\alpha}\rangle$ and $\langle\tilde{\Psi}^{tr}_{\alpha}|$, we construct the low-dimensional matrix $\begin{bmatrix}M\end{bmatrix}$ and vectors $\begin{bmatrix}\tilde{{S}}^{\dagger}\end{bmatrix}$ and $\begin{bmatrix}{S}\end{bmatrix}$—called as emulator components—at emulation stage. Note
$\begin{bmatrix}M\end{bmatrix}_{\alpha\beta}$ $=\langle\tilde{\Psi}_{\alpha}|M-M_{\beta}|\Psi_{\beta}\rangle+\langle\tilde{\Psi}_{\alpha}|S_{\beta}\rangle$ (24)
$=\langle\tilde{\Psi}_{\alpha}|M-M_{\alpha}|\Psi_{\beta}\rangle+\langle\tilde{S}_{\alpha}|\Psi_{\beta}\rangle\ .$ (25)
Since the training solution basis is not orthonormal, we need to deal with the norm matrix:
$\begin{bmatrix}N\end{bmatrix}_{\alpha\beta}$ $\equiv\langle\tilde{\Psi}_{\alpha}|\Psi_{\beta}\rangle\ .$ (26)
Now, specifically for our inhomogeneous Schrödinger equation, there, $M=E-H$ with $E$ as one important parameter. The following relations could be helpful for emulations:
$\langle\tilde{\Psi}_{\alpha}|M-M_{\beta}|\Psi_{\beta}\rangle$ $=(E-E_{\beta})\begin{bmatrix}N\end{bmatrix}_{\alpha\beta}$
$\quad-\langle\tilde{\Psi}_{\alpha}|H-H_{\beta}|\Psi_{\beta}\rangle\ .$ (27)
The projected $H$ matrix, defined as
$\begin{bmatrix}H\end{bmatrix}_{\alpha\beta}=\langle\tilde{\Psi}_{\alpha}|H|\Psi_{\beta}\rangle\ ,$ (28)
@@
$\begin{bmatrix}H\end{bmatrix}=E\begin{bmatrix}N\end{bmatrix}-\begin{bmatrix}M\end{bmatrix}\ .$ (29)
If the parameter dependencies in $\begin{bmatrix}M\end{bmatrix}$, $\begin{bmatrix}\tilde{S}^{\dagger}\end{bmatrix}$, and $\begin{bmatrix}S\end{bmatrix}$ are factorized from the high-dimensional tensors, the expensive calculations for computing the emulator components can be performed once at the training stage. Later, we scale them appropriately at the emulation stage to get those emulator components at each point with little cost. For the parameters without such affine dependence, various solutions exist, such as interpolating and extrapolating the emulator components in $\bm{\theta}$ using data-driven methods [^14] This is called emulator-in-emulator method in Ref. Zhang and Furnstahl (2022). or applying empirical interpolation methods Benner et al. (2017a) [^15] This method has been used lately, for example, in Ref. Odell et al. (2024). to approximate the non-affine dependence with affine structures. We then have fast emulations.
Note the variational approach informs the test function space—i.e., the space spanned by the $\langle\tilde{\Psi}^{tr}_{\alpha}|$—used in this work. However, other choices for the test function space have been studied in the MOR and will be explored in the future. One crucial ingredient, keeping the analyticity of $\begin{bmatrix}N\end{bmatrix}_{\alpha,\beta}$ and $\begin{bmatrix}H\end{bmatrix}_{\alpha,\beta}$ in terms of the $E^{tr}_{\alpha}$ variable, is critical for realizing non-Hermitization, which should be considered when exploring new test function space in studying quantum continuum physics.

### v2

@@ III RBM-based emulations
This section centers around the RBM emulators used to interpolate, extrapolate, or emulate $\mathcal{A}$ in a combined space of complex $E$ plane and the $\bm{\theta}$ dimensions. Note that certain parameters inside $\bm{\theta}$ could be shared among the $H$ operator and the two sources. The training points, labeled as $(E^{\mathrm{tr}}_{\alpha},\bm{\theta}^{\mathrm{tr}}_{\alpha})$ with $\alpha=1,,,N_{b}$, are those with finite (most of the time positive) $\mathrm{Im}\,E^{\mathrm{tr}}$. With those parameter values or at those points in the parameter space, I solve Eqs. (2) and (3) to get the training solution snapshots, $|\Psi^{\mathrm{tr}}_{\alpha}\rangle\equiv|\Psi(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})\rangle$ and $\langle\tilde{\Psi}^{\mathrm{tr}}_{\alpha}|\equiv\langle\tilde{\Psi}(E_{\alpha}^{\mathrm{tr}},\bm{\theta}_{\alpha}^{\mathrm{tr}})|$. This step is called the training or off-line step. In the next emulation or online step, the RBM proposes a general solution at emulation point $(E,\bm{\theta})$,
$|\Psi(E,\bm{\theta})\rangle$ $=\sum_{\alpha=1}^{N_{b}}c_{\alpha}(E,\bm{\theta})|\Psi_{\alpha}^{\mathrm{tr}}\rangle\ .$ (9)
@@
As shown later, the inhomogeneous Schrödinger equations are effectively projected into a subspace spanned by the training solutions, $|\Psi^{\mathrm{tr}}_{\alpha}\rangle$ and $\langle\tilde{\Psi}^{\mathrm{tr}}_{\alpha}|$.
@@
According to previous studies, the RBM-based emulators have robust extrapolation capabilities because the emulators learn a problem’s solution manifold—mapped out by varying the input parameters—by sampling solution vectors in the subspace [2]. Using the CEE, I extrapolate $\mathcal{A}$ at complex training energies to the real energy axis in the complex $E$ plane and further into the region where potential resonances could exist. Therefore, this procedure extracts $\mathcal{A}$ at real energies and resonance properties from the training solutions, which satisfy bound-state-like boundary conditions [18]. In addition, the CERPE emulates those continuum physics extractions in the $\bm{\theta}$ space.
@@
When emulating at real energies with $\mathrm{Im}\,(E)\to 0^{+}$, the constructed wave function in Eq. (9) has bound-state-like asymptotic behavior, which is incorrect; the correct one should be outgoing waves oscillating to infinity spatially (see Fig. 2 for a visual demonstration). How is it possible, then, for the emulated $\mathcal{A}$ at real energies to be sensible? The answer comes from the fact that the sources are spatially localized; $\mathcal{A}(E)$ can thus be determined by the overlap of the source with the internal part of the emulated solution $|\Psi\rangle$ as shown in Eq. (4). This basic setup allows me to emulate continuum physics, including $\mathcal{A}(E)$ at real $E$ and spectra (bound, resonance, and DC states), without reproducing the intricate wave function asymptotics correctly. This point will be illustrated further using an example in Fig. 2 and the discussion around it. Bear in mind that a spectrum in emulation is in its low-dimensional representation. Thus, I deem emulation as an efficient way to compress an infinite-dimensional continuous spectrum—at least a portion of it.
@@
From the discussion at the end of Sec. II.3, one can see that $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s discretized BCTs (equivalently, the DC states in the compressed spectrum) are responsible for $\begin{bmatrix}\mathcal{A}\end{bmatrix}$’s near-threshold behavior. As will be shown in the Secs. IV and V, these poles exponentially cluster toward the branch points or thresholds, similar to the pole distribution in the (near-)optimal rational approximation. The threshold locations depend only on a particular $H$, but will the distribution of the discretized BCT poles be sensitive to the sources, i.e., different dynamics? The numeric results in the Secs. IV and V suggest the answer is no, which is consistent with the distribution being the property of a spectrum (albeit compressed) but not a property unique to specific dynamics. This is also in parallel with the observation that the “widespread” behavior of the pole distribution in the (near-)optimal rational approximation of functions with branch points is determined mainly by the locations of the branch points but not significantly by the other details of the functions [96, 16].
@@
Moreover, the exponential clustering of the discretized BCT poles also means one needs a small number of them (i.e., low-dim subspace) to approximate the smooth component of $\mathcal{A}$—defined as the full $\mathcal{A}$ with the resonance contributions subtracted—over an extensive range of real energies (see e.g., Fig. 4). This explains the feasibility of compressing the DC component of the continuum spectra. Of course, the bound and resonance states are the other spectra components, but their number can not be reduced.
@@
Now, the task is to determine $c_{\alpha}(E,\bm{\theta})$’s dependence on $E$ and $\bm{\theta}$, as addressed below. As a reminder, if I fix $\bm{\theta}$ to a specific value without varying them, the emulation is called CEE; otherwise, it is named CERPE.
@@ III.1 A variational method for solving a linear system and the associated RBM ROM
Suppose I need to solve an inhomogeneous linear equation:
@@
and compute $\langle\tilde{S}|\Psi\rangle$. A variational approach exists for this task [106]. The involved functional, depending on both trial solution $|\Psi_{t}\rangle$ and an auxiliary variable $|\tilde{\Psi}_{t}\rangle$, is expressed as
@@
If a trial solution pair is close to the exact solutions, i.e., $|\Psi_{t}\rangle=|\Psi\rangle+|\delta\rangle$ and $|\tilde{\Psi}_{t}\rangle=|\tilde{\Psi}\rangle+|\tilde{\delta}\rangle$, with $|\Psi\rangle$ as Eq. (10)’s exact solution and $|\tilde{\Psi}\rangle$ as the exact solution of another equation to be defined, I will have
@@
Therefore, I require $|\tilde{\Psi}\rangle$ satisfy
[…]
so that the functional $\mathcal{F}$ is stationary around the exact solution and provides an estimate of $\langle\tilde{S}|\Psi\rangle$ on the second order of $|\delta\rangle$ and $|\tilde{\delta}\rangle$. This equation defines the adjoint state vector $|\tilde{\Psi}\rangle$, which is why $|\tilde{\Psi}\rangle$ and Eq. (3) were introduced in Sec. I.
Now, one can use the $\mathcal{F}$ functional to determine $c_{\alpha}$ as functions of $E$ and $\bm{\theta}$ in Eq. (9). This procedure leads to a choice of test functions in the framework of Galerkin projection [4]. To develop a ROM for Eq. (10) with parameters in $M$ and the sources, I construct [^12] In the previous report [1], I use the convention of the implicit sum over the repeated indices. Here, that sum is expressed explicitly.
$|\Psi_{t}\rangle$ $=\sum_{\alpha}c_{\alpha}|\Psi_{\alpha}^{\mathrm{tr}}\rangle\ ,$ (14)
$\langle\tilde{\Psi}_{t}|$ $=\sum_{\alpha}\tilde{c}_{\alpha}\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|\ ,$ (15)
@@
$M_{\alpha}|\Psi_{\alpha}^{\mathrm{tr}}\rangle$ $=|{S}_{\alpha}\rangle\ ,$ (16)
$\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|M_{\alpha}$ $=\langle\tilde{{S}}_{\alpha}|\leftrightarrow M_{\alpha}^{\dagger}|\tilde{\Psi}^{\mathrm{tr}}_{\alpha}\rangle=|\tilde{{S}}_{\alpha}\rangle\ .$ (17)
@@
[…]
$\quad-\sum_{\alpha,\beta}\tilde{c}_{\alpha}c_{\beta}\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|M|\Psi_{\beta}^{\mathrm{tr}}\rangle\ .$ (18)
Note when S{S}, $\tilde{{S}}$, or $M$ is without $\alpha.\beta$ indices, they are evaluated at the general emulation point. To find the stationary points, I look for the solutions such that the first derivatives of $\mathcal{F}$ with respect to $c$ and $\tilde{c}$ are zero. One then obtains a low-dimensional linear equation system:
$\begin{pmatrix}0&\begin{bmatrix}M\end{bmatrix}^{T}\\ \begin{bmatrix}M\end{bmatrix}&0\end{pmatrix}\begin{pmatrix}c\\ \tilde{c}\end{pmatrix}=\begin{pmatrix}\begin{bmatrix}\tilde{S}^{\dagger}\end{bmatrix}\\ \begin{bmatrix}S\end{bmatrix}\end{pmatrix}\$. (19)
The upper right corner of the matrix on the left side uses transpose, T, not complex conjugate, due to my definition of $\tilde{c}$ in Eq. (14). The block matrix has its matrix elements defined as
$\begin{bmatrix}M\end{bmatrix}_{\alpha,\beta}\equiv\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|M|\Psi_{\beta}^{\mathrm{tr}}\rangle\,$, (20)
and similarly, the projected source vectors have
$\begin{bmatrix}\tilde{{S}}^{\dagger}\end{bmatrix}_{\alpha}$ $\equiv\langle\tilde{{S}}|\Psi_{\alpha}^{\mathrm{tr}}\rangle\,$, (21)
$\begin{bmatrix}{S}\end{bmatrix}_{\alpha}$ $\equiv\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|{S}\rangle\ .$ (22)
@@
$\sum_{\alpha}c_{\alpha}\langle\tilde{{S}}|\Psi_{\alpha}^{\mathrm{tr}}\rangle$ $=\sum_{\alpha}\tilde{c}_{\alpha}\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|{S}\rangle$
$=\sum_{\alpha,\beta}\tilde{c}_{\alpha}c_{\beta}\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|M|\Psi_{\beta}^{\mathrm{tr}}\rangle\ .$ (23)
One thus has
$\mathcal{F}|_{\mathrm{stationary}}$ $=\sum_{\alpha}c_{\alpha}\langle\tilde{S}|\Psi_{\alpha}^{\mathrm{tr}}\rangle$ (24)
Therefore, after obtaining, $|\Psi^{\mathrm{tr}}_{\alpha}\rangle$ and $\langle\tilde{\Psi}^{\mathrm{tr}}_{\alpha}|$, I construct the low-dimensional matrix $\begin{bmatrix}M\end{bmatrix}$ and vectors $\begin{bmatrix}\tilde{{S}}^{\dagger}\end{bmatrix}$ and $\begin{bmatrix}{S}\end{bmatrix}$—called as emulator components—at emulation stage. Note
$\begin{bmatrix}M\end{bmatrix}_{\alpha\beta}$ $=\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|M-M_{\beta}|\Psi_{\beta}^{\mathrm{tr}}\rangle+\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|S_{\beta}\rangle$ (25)
$=\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|M-M_{\alpha}|\Psi_{\beta}^{\mathrm{tr}}\rangle+\langle\tilde{S}_{\alpha}|\Psi_{\beta}^{\mathrm{tr}}\rangle\ .$ (26)
Since the training solution basis is not orthonormal, I need to deal with the norm matrix:
$\begin{bmatrix}N\end{bmatrix}_{\alpha\beta}$ $\equiv\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|\Psi_{\beta}^{\mathrm{tr}}\rangle\ .$ (27)
Now, specifically for the inhomogeneous Schrödinger equation, there, M=EHM=E-H with $E$ as one important parameter. The projected $H$ matrix, defined as
$\begin{bmatrix}H\end{bmatrix}_{\alpha\beta}=\langle\tilde{\Psi}_{\alpha}^{\mathrm{tr}}|H|\Psi_{\beta}^{\mathrm{tr}}\rangle\ ,$ (28)
@@
$\begin{bmatrix}H\end{bmatrix}=E\begin{bmatrix}N\end{bmatrix}-\begin{bmatrix}M\end{bmatrix}\$. (29)
If the parameter dependencies in $\begin{bmatrix}M\end{bmatrix}$, $\begin{bmatrix}\tilde{S}^{\dagger}\end{bmatrix}$, and $\begin{bmatrix}S\end{bmatrix}$ are factorized from the high-dimensional tensors, the expensive calculations for computing the emulator components can be performed once at the training stage. Later, one can scale them appropriately at the emulation stage to get those emulator components at each point with little cost. For the parameters without such affine dependence, various solutions exist, such as interpolating and extrapolating the emulator components in $\bm{\theta}$ using data-driven methods (e.g. the emulator-in-emulator method in Ref. [40]) or applying empirical interpolation methods [21, 45] to approximate the non-affine dependence with affine structures. One then have fast emulations.
Note the variational approach informs the test function space—i.e., the space spanned by the $\langle\tilde{\Psi}^{\mathrm{tr}}_{\alpha}|$—used in this work. However, other choices for the test function space have been studied in the MOR and will be explored in the future. One crucial ingredient, keeping the analyticity of $\begin{bmatrix}N\end{bmatrix}_{\alpha,\beta}$ and $\begin{bmatrix}H\end{bmatrix}_{\alpha,\beta}$ in terms of the $E^{\mathrm{tr}}_{\alpha}$ variable, is critical for realizing non-Hermitization, which should be considered when exploring new test function space in studying quantum continuum physics.